OpenStudy (anonymous):

How many different 5 card hands can be drawn from a standard deck of 52 cards?

Parth (parthkohli):

52C5

OpenStudy (lgbasallote):

i believe you use combination \[\huge \frac{n!}{r!(n-r)!}\]

Parth (parthkohli):

You can simplify that thing: \(\Rightarrow \Large {52! \over (52 - 5)! 5!} \)

OpenStudy (lgbasallote):

where n = 52 and r = 5

OpenStudy (lgbasallote):

got it @shivraj

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

great! glad to hear :D

OpenStudy (anonymous):

\[^nC_r = \frac{n!}{(n-r)! \times r!}\]

Parth (parthkohli):

\(\Rightarrow \Large {52 \times 51\times 50 \times 49 \times 48 \over 5 \times 4 \times 3 \times 2 \times 1 }\)